There are 2 types of popcorn and 4 types of hotdogs available at a concession stand in a cinema.
1. How many types of food are available at the stand?
2. Aside from the food above, the stand also provides 5 types of drinks. Eddie has a coupon that gives a 10% off the price for buying one type of food together with with a drinks. If eddie is going to use the coupon, how many choices does he have?

Found 2 solutions by ikleyn, ewatrrr:
Answer by ikleyn(52778)

.

(1)  2 + 4 = 6.      ANSWER



(2)  2*5 + 4*5 = 10 + 20 = 30.      ANSWER



Answer by ewatrrr(24785)


Hi
1. 2+4 = 6   types of food are available at the stand
2. Five types of beverages
 Choices of a Food/beverage:  (6C1)(5C1) 
Wish You the Best in your Studies.